OpenStudy (anonymous):

What is the formula for Maximum range and Maximum height? Medals will be rewarded.

OpenStudy (paxpolaris):

range & height of what? be specific

OpenStudy (anonymous):

in projectile motion

OpenStudy (paxpolaris):

as the projectile goes upwards it slows down ... til it's vertical velocity becomes 0. Then it falls back down to find max height: you can use the formula involving the squares of the velocities

OpenStudy (paxpolaris):

\[\LARGE {v_y}^2={u_y}^2-2g\cdot y\] where, \(y\) is the vertical position or height you are solving for (assuming you throw the projectile from y=0). \(u_y\) is the vertical component of the initial/ starting velocity. \(v_y\) is the vertical velocity at the point you trying to calculate the height for. (at max height \(v_y=0\)) \(g\) is acceleration due to gravity \((\approx9.8m/s^2\) or \(32ft/s^2)\) \[\large \implies 0 = {u_y}^2 -2g \cdot y_\max\]\[\large \implies y_\max = {{u_y}^2 \over2g }\]

OpenStudy (paxpolaris):

to find max range: .... assuming the initial velocity is constant, only the angle you throw the projectile can change. the max range is when throwing at a 45 degree angle. \[\Large x = {2u^2 \over g}\cdot \sin(\theta)\cdot \cos(\theta)\]\[\large \implies x_\max ={u^2 \over g} \ \dots \text{{at 45 degrees}}\] where, \(x\) is range: the horizontal position where the projectile falls back on the ground \(u\) is initial velocity